OpenStudy (anonymous):

(³√x^2) * (√x) I made up this equation to show as a an example for a problem, and got to here: x^2/x^3 * (x/x^2) is there anything more I can do?

OpenStudy (turingtest):

use\[\large\sqrt[b]{x^a}=x^{a/b}\]and\[\large x^a\cdot x^b=x^{a+b}\]

OpenStudy (turingtest):

what you have written is not correct I'll walk you through this one since it is just a practice problem you say you made up...

OpenStudy (turingtest):

\[\large\sqrt[3]{x^2}\cdot\sqrt x=x^{2/3}\cdot x^{1/2}=x^{\frac23+\frac12}=x^{\frac46+\frac36}=x^{7/6}=\sqrt[6]{x^7}\]

OpenStudy (turingtest):

make sense?

OpenStudy (anonymous):

Yes, thank you very much

OpenStudy (turingtest):

welcome!